OpenStudy (anonymous):

WILL MEDAL & FAN. Solve for r, I have the answer but im confused on how it was gotten P= A/1+rt do I multiply by the denominator? Answer is: r= A-P/pt Posting this again cause I cant get anyone to help and this bump thing doesn’t seem to be working…

OpenStudy (litening1337):

5. YW.

OpenStudy (anonymous):

there are no numbers?

OpenStudy (anonymous):

P=A/(1+rt) Multiply both sides by (1+rt) P(1+rt)=A Distribute P P+Prt=A Subtract P from both sides Prt=A-P Divide both sides by Pt r=(A-P)/(Pt)